OpenStudy (anonymous):

Find the area of a sector with a central angle of 216 degrees and a radius of 18.4 millimeters and round to the nearest tenth. Help please!

OpenStudy (anonymous):

do u know formula \[A=\frac{ 1 }{ 2 } r ^{2} \theta \]

OpenStudy (anonymous):

No. I missed a day of class and that was the only day my teacher went over that and I have an assignment on it. :(

OpenStudy (anonymous):

dont use the formula i gave u its in radians

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

area =\[\pi r ^{2} (\frac{ c }{ 360 })\]

OpenStudy (anonymous):

c is the central angle in ur case its 216

OpenStudy (anonymous):

\[\pi 18.4^{2} \times \frac{ 216 }{ 360 } =638.1705653 mm ^{2}\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

So that's the formula I should use?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome, thank you so much!

OpenStudy (anonymous):

\[area= \pi r ^{2}\times \frac{ c }{ 360 }\]

OpenStudy (anonymous):

no worries :)